OpenStudy (anonymous):

find the equation of the line tangent to y=2secx at =pi/4 , someone can plz explain me step by step??

OpenStudy (jdoe0001):

tangent to y=2secx at x=pi/4 ?

OpenStudy (jdoe0001):

someone, no name, just deleted 2 lines :P

OpenStudy (anonymous):

oohh yes its x=pi/4

OpenStudy (jdoe0001):

keep in mind that a derivative is the EQUATION TO FIND THE SLOPE for that function is not the slope value itself so what would be the derivative of 2sec(x)?

OpenStudy (anonymous):

2secxtanx

OpenStudy (whpalmer4):

Okay, so the slope of tangent line at \(x=\frac{\pi}{4}\) is the value of the first derivative at that same value of \(x\): \[m = 2\sec (\frac{\pi}{4})\tan(\frac{\pi}{4}) =\] Now use that in addition to the point-slope formula to establish the tangent line: \[y-y_0 = m(x-x_0)\]You'll need to evaluate \(y_0 = 2\sec(x_0)\) with \(x_0 = \dfrac{\pi}{4}\) to fill in all the blanks.

OpenStudy (whpalmer4):

If you do it correctly, you'll end up with a tangent line like the one illustrated in the attached graph.

OpenStudy (anonymous):

do i have to draw that ?

OpenStudy (whpalmer4):

I don't know, does your problem ask you to do so? I'm just illustrating the result. If you made a mistake, you might end up with something like one of these two:

OpenStudy (whpalmer4):

I'm a big fan of checking answers. A graph is an obvious way to check the answer to a problem like this, wouldn't you say?